give the usual ax^2 + bx + c form of the quadratic function wich has a= 1, and has two zeros z= 3 + 2i and z'= 3 - 2i
Answer by edjones(8007)

x=3+2i, x=3-2i
=(x-(3+2i)) (x-(3-2i)
=x^2-6x+13
.
Ed